OpenStudy (anonymous):

1) What are all the 4 digit numbers you can write using the digits 2, 5, 7, and 9? Each digit only once in a number. please write the possible numbers. 2) Rocco has five 20 cent stamps and 3 35 cent stamps. What amounts of postage can he make using one or more stamps? please show all possible ways and show work.

OpenStudy (anonymous):

1) you could make 4! = 4*3*2*1 = 24 four digits numbers with 2, 5, 7, and 9 2) in this one you have to find all possible combinations of 1 stamp, 2 stamps, 3 stamps ..... 8 stamps

OpenStudy (anonymous):

@swissgirl

OpenStudy (swissgirl):

wait ill brb

OpenStudy (anonymous):

k

OpenStudy (swissgirl):

1) ok we have 4 digits, 2,5,7and 9 and we would like to know all the possible combinations So the first place we have 4 options, in the second place we have 3 options since we chose our first option, in third place we have 2 options since we have chosen our first 2 options and in our last place we have one option cuz the other 3 digits have been placed already so that is 4*3*2*1=24 So we there are 24 possible combinations of these 4 digits

OpenStudy (anonymous):

2

OpenStudy (swissgirl):

waittt a sec I unsure what the first part of the question is actually saying. I am not sure if it wants the amount or the diff options I guess you can state both 2579 5279 7529 9572 2597 5297 7592 9527 2957 5927 7952 9725 2975 5972 7925 9752 2795 5792 7295 9275 2759 5729 7259 9257 Here are all the possible combinations and if u will count them there will be 24 diff combinations

OpenStudy (swissgirl):

This is still for part 1

OpenStudy (anonymous):

could you please help me with # 2?

OpenStudy (swissgirl):

yes give me a few seconds

OpenStudy (swissgirl):

just have to write out all the options Do u need a formula or do u just have to list all teh options

OpenStudy (anonymous):

please

OpenStudy (swissgirl):

Do u need a formula or u need to write out all the options or u just dont get what i am even saying

OpenStudy (anonymous):

formula

OpenStudy (swissgirl):

ohhhhhhh okkk that is simpler

OpenStudy (anonymous):

or you could give me the options if you dont know the formula

OpenStudy (swissgirl):

nah i can figure out the formula i think its simple

OpenStudy (anonymous):

k

OpenStudy (swissgirl):

There can be repeats and I dont know a formula for repeats hmmmmmm

OpenStudy (anonymous):

there might not be a formula

OpenStudy (swissgirl):

Still working on it give me a few more seconds

OpenStudy (anonymous):

k

OpenStudy (swissgirl):

ok i realized a pattern already but i didnt finish writing out all the options 2+3+4+5+6+7+8+9=44 there shld be 44 options Like when we have one stamp there are 2 options When we have 2 stamps there are 3 options and so forth but I didnt finish getting all the options

OpenStudy (swissgirl):

20 35 20 20 35 35 20 35 20 20 20 20 35 20 35 20 25 35 35 35 20 20 20 20 20 35 35 35 20 20 35 35 20 20 20 35 35 35 35 35 20 20 20 20 20 20 35 35 35 35 20 20 35 35 35 20 20 20 35 35 20 20 20 20 35 35 35 35 35 35 20 20 20 20 20 20 20 35 35 35 35 35 20 20 35 35 35 35 20 20 20 35 35 35 20 20 20 20 35 35 20 20 20 20 20 35 35 35 35 35 35 35 20 20 20 20 20 20 20 20 35 35 35 35 35 35 20 20 35 35 35 35 35 20 20 20 35 35 35 35 20 20 20 20 35 35 35 20 20 20 20 20 35 35 20 20 20 20 20 20 35 35 35 35 35 35 35 35 20 20 20 20 20 20 20 20 20 35 35 35 35 35 35 35 20 20 35 35 35 35 35 35 20 20 20 35 35 35 35 35 20 20 20 20 35 35 35 35 20 20 20 20 20 35 35 35 20 20 20 20 20 20 35 35 20 20 20 20 20 20 20 35 35 35 35 35 35 35 35 35 There are the 44 options

OpenStudy (swissgirl):

hahaah I dont know of a formula but all i noticed was that each option had one more so option #1 =2 option#2= option#1 +1 option #3= option#2 +1 option #4 = option#3 +1 So its kinda like a recurring sequence that was the best formula i thought of lol

OpenStudy (anonymous):

k thanks

OpenStudy (swissgirl):

is this what u needed?

OpenStudy (anonymous):

yes

OpenStudy (swissgirl):

Great :)